### v1

# Analyzing and Mitigating Bias for Vulnerable Classes: Towards Balanced Representation in Dataset

Abstract
The accuracy and fairness of perception systems in autonomous driving are crucial, particularly for vulnerable road users. Mainstream research has looked into improving the performance metrics for classification accuracy. However, the hidden traits of bias inheritance in the AI models, class imbalances and disparities in the datasets are often overlooked. In this context, our study examines the class imbalances for vulnerable road users by focusing on class distribution analysis, performance evaluation, and bias impact assessment. We identify the concern of imbalances in class representation, leading to potential biases in detection accuracy. Utilizing popular CNN models and Vision Transformers (ViTs) with the nuScenes dataset, our performance evaluation reveals detection disparities for underrepresented classes. We propose a methodology for model optimization and bias mitigation, which includes data augmentation, resampling, and metric-specific learning. Using the proposed mitigation approaches, we see improvement in IoU(%) and NDS(%) metrics from 71.3 to 75.6 and 80.6 to 83.7 respectively, for the CNN model. Similarly, for ViT, we observe improvement in IoU and NDS metrics from 74.9 to 79.2 and 83.8 to 87.1 respectively. This research contributes to developing more reliable models and datasets, enhancing inclusiveness for minority classes.
I Introduction
The advent of autonomous driving technology marks a significant leap forward in the automotive industry. While the benefits of autonomous vehicles (AVs) are manifold, including potential reductions in traffic accidents and increased mobility for various demographics, the technology’s reliability is highly contingent on the accuracy and fairness of its perception systems [12] [14]. These systems, particularly those involving object detection, are critical for identifying and appropriately reacting to various entities on the road, such as pedestrians, cyclists, motorcyclists, and mobility-impaired persons [8] [31]. However, recent studies have underscored the existence of biases in autonomous driving datasets, which could lead to discriminatory treatment of certain road users [17] [25]. Our study delves into two prominent datasets in this domain, nuScenes [4] and Waymo [27], aiming to uncover and address these biases. We explore the representation of vulnerable classes on object detection models’ accuracy and fairness through class distribution analysis, model’s performance evaluation, and bias impact assessment. Class imbalance refers to the disproportionate representation of different categories or classes within a dataset [23]. In the context of autonomous driving, this often manifests as an overrepresentation of certain types of vehicles, while pedestrians, cyclists, motorcyclists, and especially mobility-impaired individuals are underrepresented [17]. The imbalance is not only a statistical concern but a practical one, as it can lead to perception systems that are less adept at recognizing minor, yet crucial classes. Studies have shown that this imbalance can result in biased learning outcomes, where models preferentially detect overrepresented classes [29]. The consequences of such an imbalance in the dataset cannot be overlooked, as evidenced by incidents like the Uber self-driving car collision [26]. In this dataset imbalance context, our paper explores:
  1. 1.
    How do the representation and disparities of classes in datasets impact the accuracy of AI models?
  2. 2.
    How do biases in AI systems for detecting under-represented classes manifest in visual application, and what implications do they have on diverse models?
To counter biases towards learning, our research explores a framework for balanced class representation and AI model optimization. We emphasize the need for diverse and representative datasets that include a wide range of scenarios, reflecting the complexity of real-world conditions. The proposed framework includes bias assessment and model re-calibration to ensure model fairness and effectiveness with the retraining of models with the inclusion of new classes/sub-samples. We also highlight critical gaps in current vision datasets used in autonomous driving through metric-specific learning to develop more equitable and reliable systems. The aim is to contribute to the generalization of AI algorithms and respective datasets.
II Background and Related Work
Perception technology in autonomous driving has evolved through extensive research and development, with significant milestones achieved in sensor technology, algorithmic processing, and machine learning models. Early systems relied heavily on rule-based algorithms and limited sensor input, but the advancements in deep learning and improvements in LiDAR, radar, and camera technologies have provided next-generation solutions [15]. Despite these advancements, the challenge of developing perception systems that can reliably interpret diverse and unpredictable road scenarios remains a significant research focus. Datasets play a crucial role in training and validating autonomous driving systems. Datasets like KITTI [9] and Cityscapes [6] have provided foundations for advancing the field. More recently, datasets such as nuScenes and Waymo have offered more diverse and complex environments for training and testing AI models. These datasets include various annotations for vehicles, pedestrians, cyclists, and other entities, providing a rich foundation for developing and evaluating perception algorithms.
Class Imbalance: Efforts to address class imbalance and bias in autonomous driving datasets have included techniques like data augmentation, re-sampling, and advanced algorithmic approaches like cost-sensitive learning [11]. Research in this area seeks to balance the representation of classes in training data [28] and develop fair and accurate models across all categories [29] [31]. For instance, in research paper[10], the author highlights the use of scenario simulations to test and improve model performance in detecting imbalanced classes. The ethical implications of biased autonomous driving systems are significant. There is a growing discourse on the need for equitable AI systems that ensure the safety and fairness of all road users [24]. Furthermore, regulatory bodies are increasingly focusing on how these technologies adhere to safety standards, particularly in terms of their ability to detect and respond to diverse and representative [5] scenarios. Targeting imbalanced representation, a novel approach that combines visual codebook generation with deep features and a non-linear $Chi^{2}$ SVM classifier is porposed in [23]. This method tackles the issue of imbalanced datasets, where traditional algorithms often fail to detect minority classes effectively. The approach involves extracting low-level deep features using transfer learning with the ResNet-50 pre-trained network and k-means clustering to create a visual codebook. Each image is then represented as a Bag-of-Visual-Words (BOVW), derived from the histogram of visual words in the vocabulary. The $Chi^{2}$ SVM classifier is used for classifying these features, demonstrating optimal performance in empirical analysis. This method shows superior accuracy, F1-score, and AUC metrics results compared to state-of-the-art methods, as validated on two challenging datasets: Graz-02 and TF-Flowers [23].
With a focus on training methods to address the challenges of learning from imbalanced data, a new loss function that mitigates the impact of samples leading to overfitted decision boundaries is proposed in [21]. This loss function has been shown to enhance the performance of various imbalance learning methods. The proposed approach is versatile and can be integrated with existing resampling, meta-learning, and cost-sensitive learning methods to tackle class imbalance problems. The paper validates this method with experiments on multiple benchmark datasets, demonstrating its superiority over state-of-the-art cost-sensitive loss methods. Additionally, the authors have made their code publicly available, facilitating further research and application [21]. A novel approach for predicting driver behaviour, which is crucial for safely integrating autonomous vehicles into human-dominated traffic, is presented in [3]. The proposed method addresses the shortcomings of existing predictive models, which either lack transparency (deep neural networks) or are not expressive enough (rule-based models). The authors introduce a model that embeds the Intelligent Driver Model (IDM), a rule-based approach, into deep neural networks. This hybrid model combines the long-term coherence and interpretability of rule-based models with the expressiveness of deep learning, aiming to accurately predict driver behaviour in complex scenarios like merging. The method is an attempt to bridge the gap between two modelling paradigms. It enhances the interpretability of neural network predictions while maintaining accuracy, a critical factor for real-time decision-making in autonomous driving. The model’s transparency is particularly beneficial for debugging and understanding its predictions.
Model learning and Representation: The paper by Wang et al. [30] introduces a deep attention-based method for imbalanced image classification (DAIIC), utilizing an attention mechanism within a logistic regression framework to prioritize minority classes in prediction and feature representation. It automatically determines misclassification costs to aid discriminative feature learning. Robust across different networks and datasets, the DAIIC method has proven effective, surpassing several benchmarks in single-label and multi-label contexts [30]. The background and related work in autonomous driving technologies, datasets, and the challenges of class imbalance and bias underscore the importance of this research area. Our study builds upon these foundations, aiming to contribute to the development of more equitable and reliable autonomous driving systems. By addressing class imbalance and bias in key datasets, we seek to enhance the safety and fairness of these technologies for all road users.
III Methodology
When an imbalanced dataset $D$ with an unjustified distribution of classes ${C_{1},C_{2},.....,C_{n}}$ is used to train machine learning models, specifically neural networks and Vision Transformers (ViTs), the models are likely to inherit biases because of class representation in the dataset. These biases may arise due to imbalanced class representation and a predominance of normal conditions as compared to challenging real-world scenarios. For neural networks, we anticipate biases to be reflected in neurons’ sensitivity and selectivity scores, while for vision transformers, biases may manifest in the attention allocation. Therefore, we propose a methodology consisting of statistical analysis for analyzing and mitigating biases using dataset analysis, bias impact assessment, bias mitigation, metrics behaviour analysis, and bias assessment and model re-calibration, as shown in Figure 1. The following subsections detail and discuss the concepts and terminologies used in our methods and tests.
Figure 1: Proposed methodology for analyzing and mitigating biases
III-A Dataset and Classes
In this paper, we use nuScenes as the case study, focusing on the Pedestrian, Cyclist, and Motorcyclist classes. However, the approach can be tested and validated on other driving datasets, e.g., the Waymo dataset [27], where the samples can be Vehicles, Pedestrians, and Cyclists.
III-A1 3.1.1 Class Distribution Analysis
We perform statistical analysis on the frequency of the specified classes in the nuScenes and Waymo datasets to identify class imbalances. The datasets are scrutinized to understand the context (urban versus rural settings, diverse weather and lighting conditions) in which these classes are represented, highlighting any underrepresented scenarios.
Pedestrian Class: nuScenes dataset includes a significant proportion of pedestrian annotations, with 149,921 instances labelled as “human.pedestrian.adult” accounting for 21.61% of all annotations. This substantial representation highlights the importance of pedestrian detection in autonomous driving systems. However, other pedestrian sub-categories like children (1,934 annotations, 0.28%), construction workers (13,582 annotations, 1.96%), and individuals with personal mobility devices (2,281 annotations, 0.33%) are less represented. These differences underscore potential gaps in the dataset’s and subsamples diversity.
Cyclist Class: It is categorized as “vehicle.bicycle” and has around 17,060 annotations, comprising 2.46% of the total dataset. This relatively lower representation compared to pedestrians impacts the class weight calculation and therefore the model’s ability to accurately detect cyclists, which might become a critical safety concern in urban driving scenarios.
Motorcyclist Class: It is annotated as “vehicle.motorcycle” and has around 16,779 annotations, making up 2.42% of the dataset. This class is slightly underrepresented compared to pedestrians but is on par with cyclists. The lesser representation of cyclists and motorcyclists indicates an area for improvement in class balance, especially considering the different dynamics and associated risks.
The common metrics used for model performance evaluation in nuScenes are, mean average Precision (mAP), Intersection over Union (IoU), and nuScenes Detection Score (NDS). mAP and IoU are well-known metrics used in 3d object detection and segmentation tasks [4], and are calculated using the fundamentals of precision, recall, area of overlap and intersection over union respectively. Similarly, NDS combines several metrics, including the mAP, to provide a comprehensive performance score for the object detection models. The NDS metric is calculated as:
$NDS=\frac{1}{10}[5\cdot mAP+\sum_{mTP\in TP}(1-\min(1,mTP))]$ (1)
III-A2 3.1.2 Data Augmentation/Re-sampling
Given the class representation imbalances, particularly in the less represented sub-categories of pedestrians and the overall lower representation of cyclists and motorcyclists, specific strategies can be employed for model training and evaluation.
Resampling Technique: Addressing class imbalances in the datasets, particularly for underrepresented classes like cyclists and motorcyclists, is crucial for unbiased model training in autonomous driving systems. We apply both “Random Oversampling” and “Undersampling ” from resampling techniques”. Random Oversampling duplicates instances from minority classes, ensuring models have sufficient data to learn from these critical yet less frequent road users. Conversely, Undersampling reduces the predominance of over-represented classes, like pedestrians, to prevent bias inheritance in models. Both methods are carefully balanced to maintain data diversity [20].
Data Augmentation: Another key approach is the use of geometric transformations, such as rotating and flipping images. This technique represents various class orientations, which is crucial for autonomous driving where objects are encountered from multiple directions. These transformations not only prevent the model from developing orientation-specific biases but also enhance its ability to adapt to diverse real-world scenarios. Additionally, this approach contributes to increasing the dataset’s diversity and size, thereby improving the model’s robustness and generalization capabilities for more accurate and reliable perception tasks [22].
Combination of Strategies: To further have varied training and testing and to increase the samples/sub-classes diversity, we combine resampling and data augmentation to balance the training dataset. This approach integrates the strengths of both Resampling (Random undersampling/Oversampling) and Transformations, such as rotating and flipping images, to optimize the representation and variability of data.
III-B Models and Behavioural Metrics
We strategically use popular models such as ResNet18, SqueezeNet from the CNN and ViT from the transformer family for evaluation. The selection of these models is intended to provide an understanding of how different architectures perform on biased datasets and to compare their ability to generalize across various classes, especially after the implementation of bias mitigation. Using these models, we further gain insights across different neural network architectures, ensuring robustness and reliability in our findings.
Model Selection: ResNet18 [18] is a well-known architecture and performs fairly in perception tasks due to its ability to capture spatial hierarchies, but it can show a tendency of bias towards common textures and patterns. SqueezeNet [13], is a compact model and focuses on global image features but risks inheriting biases from non-diverse datasets [14]. Vision Transformers (ViTs) process images by analyzing sequences of patches, adeptly focusing on relevant image parts due to their attention mechanisms [7]. While ViTs are proficient at capturing dependencies, they may struggle to adequately focus on underrepresented classes/unique scenarios. Each model’s inherent strengths and learning mechanisms influence classification and learning, underscoring the need for tailored bias detection and mitigation strategies.
Behaviour Metrics: Sensitivity and selectivity scores of neurons are used to evaluate biases in models like ResNet and SqueezeNet. Sensitivity scores measure a neuron’s response to changes in input from specific classes, indicating potential biases. For example, a ResNet neuron more sensitive to vehicles than cyclists might show a vehicle-detection bias. Selectivity scores, on the other hand, assess the specificity of a neuron’s response to a class, with high selectivity indicating specialized recognition capabilities. This helps identify whether models effectively differentiate between classes or exhibit biases, such as a score showing higher selectivity for pedestrians over motorcyclists, potentially indicating a feature recognition bias. The formula is described as:
Sensitivity $\text{Sensitivity Score}=\partial a/\partial x\text{\;}$
where $a$ is activation of neuron, and $x$ is the input feature.
Selectivity $\text{Selectivity Score}=\frac{a_{c}-a_{avg}}{\max(a_{c},a_{avg})}\text{\;}$
where $a_{c}$ is activation of the neuron for the target class and $a_{avg}$ is the average activation for all classes.
In Vision Transformers (ViTs), attention map analysis is key for identifying model focus areas and potential decision-making biases. This involves extracting attention weights from each Transformer layer, which indicates the model’s focus on different image parts. These heatmaps created using the scaled dot-product attention mechanism, help determine if the model disproportionately focuses on certain features, potentially signaling bias. Attention weights are calculated as $A=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$, where $Q$ (Query), $K$ (Key), and $V$ (Value) are matrices derived from the input, and $d_{k}$ is the dimension of the Key vectors. By analyzing these attention maps across different layers and attention heads, we can identify whether the model disproportionately focuses on certain features or misses critical aspects, potentially indicating bias.
Layer-wise Relevance Propagation (LRP) offers another approach, starting from the output layer and backpropagating the relevance score of the predicted class through the network. LRP redistributes each layer’s relevance to its inputs, particularly in self-attention layers, where relevance among patches is allocated based on attention weights [1]. The relevance for a patch in a layer, highlighting the model’s decision-making basis is calculated as:
$R_{j}^{(l)}=\sum_{i}A_{ij}^{(l)}R_{i}^{(l+1)}$ (2)
Where $R_{j}^{(l)}$ is the relevance of patch $j$ in layer $l$, $A_{ij}^{(l)}$ are the attention weights from patch $j$ to patch $i$ in the self-attention mechanism, and $R_{i}^{(l+1)}$ is the relevance in the next layer.
III-C Bias Impact Assessment
This section describes model errors (false positives and negatives) and examines learning patterns and bias inheritance, crucial for identifying potential biases in model performance.
III-C1 3.3.1 Error Analysis
We focus on false positives (e.g., incorrectly identifying objects as belonging to a target class) and false negatives (e.g., failing to detect cyclists). We further use Class-specific error analysis, which includes evaluating how well the model detects pedestrians, cyclists, and motorcyclists under various scenarios. We also examine if there is a tendency to overlook certain classes and how this impacts overall model performance. The correlation between these errors and previously identified biases, such as neurons’ sensitivity or selectivity scores, is also explored to understand the underlying causes and guide bias mitigation strategies using bias correlation.
III-C2 3.3.2 Learning Patterns and Bias Inheritance
Different model architectures lead to varied learning behaviours and potential biases. CNNs like ResNet may focus on textures and local patterns, which could introduce biases if such features are not uniformly present across all classes. This can result in misclassification, especially when distinguishing features are missing. In contrast, Vision Transformers (ViTs) may develop biases based on how attention is distributed across an image.
Bias Inheritance from Data: A model’s performance can vary depending on the diversity/samples encountered in training. For instance, a model trained predominantly on urban pedestrian images might underperform in rural settings. Understanding this bias inheritance is crucial for ensuring that perception systems are adaptable to diverse real-world conditions and can guide necessary adjustments in data representation and model training.
Comparative Analysis: By comparing how different models like ResNet, SqueezeNet, and ViTs respond to the same dataset, we can uncover specific biases inherent in each architecture. This comparative approach helps identify which models are more prone to biases, helping categorise the most suitable model for bias-sensitive applications.
Interpretation and Visualization: Layer-wise Relevance Propagation (LRP) provides insights into what drives a model’s decisions by highlighting influential input features. For example, if a model consistently focuses on irrelevant background features for decision-making, it indicates a bias that needs correction [19]. Analysis of these scores, using visual tools like heatmaps and attention maps, helps in interpreting these complex models, offering an understanding of their decision-making processes, and adjusting the training process to mitigate identified biases.
Overall, this process involving analysis of learning patterns, data bias inheritance comparative model analysis, and visualization can help discover bias occurrence/inheritance in different model architectures.
III-D Bias Mitigation Techniques
To mitigate biases inherited from sample representation, we explore the strategies mentioned below to address these class disparities. These techniques are crucial for addressing the disparities, ensuring that the model’s performance is equitable and reliable across all categories.
III-D1 3.4.1 Cost-Sensitive Learning
Customized cost-sensitive learning offers a robust approach to mitigating these issues by adjusting the model’s loss function according to the representation of each class. The class weights calculation and loss function are discussed below.
Class Weights Calculation: Based on the dataset statistics, we calculate the weights for the pedestrian ${w_{p}}$, cyclist ${w_{c}}$, and motorcyclist ${w_{m}}$ classes using the inverse of their respective representation percentages. This approach assigns higher weights to underrepresented classes, emphasizing their importance during training. The calculated weights are:
[…]
These weights are then normalized to ensure they contribute proportionally and maintain stability during training.
Loss Function: We integrate these calculated weights into the model’s loss function, utilizing a weighted multi-class cross-entropy loss. For each instance i, the loss is defined:
[…]
Here ${w_{p}}$, ${w_{c}}$, and ${w_{m}}$ are the weights for pedestrians, cyclists, and motorcyclists. ${y_{ix}}$ is a binary indicator of whether instance i belongs to a class x. ${p_{ix}}$ is the model’s predicted probability for instance i belongs to a class x.
Dynamic Adjustment Evaluation: To ensure continual model adaptability we propose dynamically adjusting these weights based on performance metrics. This approach can be tested using a validation set to monitor performance improvements for underrepresented classes and to prevent overfitting.
By employing this customized cost-sensitive learning approach, we aim to reduce the biases inherent in the AI model from the imbalanced dataset, ensuring a more equitable and robust model performance across all classes, which is crucial for the reliability and safety of perception applications.
III-D2 3.4.2 Data Augmentation Based on Model Analysis
We use a targeted data augmentation strategy to address biases identified in ViTs through attention map analysis and Layer-wise Relevance Propagation (LRP). This approach specifically addresses under-representation issues in cyclists and motorcyclists within the nuScenes dataset.
Attention-Guided Augmentation: Insights from attention maps guide this augmentation strategy. For instance, if the model frequently overlooks pedestrians under night or certain lighting conditions, then there is a need for more class samples highlighting these scenarios. Techniques such as zooming or adjusting brightness/contrast are used to emphasize these aspects. Similarly, additional variations of poses or orientations for motorcyclists are incorporated to improve model recognition abilities in these contexts.
LRP-Informed Sampling: Layer-wise Relevance Propagation provides a deeper understanding of which features contribute most significantly to the model’s decisions. Using insights from LRP, we can selectively augment our dataset to enhance the representation of features critical for correct classifications. Augmentation for this case includes scenarios where the model typically misclassifies images, like partially occluded subjects or specific textures and patterns, enhancing the model’s ability to differentiate between relevant and misleading features.
Implementation:This strategy involves iterative dataset refinement based on continuous model analysis. By aligning augmentation closely with model needs and performance, we aim to mitigate biases, effectively ensuring balanced and fair model performance. This is critical for perception systems, where the accurate detection of all road users, particularly those underrepresented like cyclists and motorcyclists.
III-E Bias Assessment and Model Re-calibration
Effective bias mitigation requires continuous evaluation and adjustment of models. This section describes our approach to bias assessment and model re-calibration.
Bias Assessment: In this step, we use behaviour metrics, specifically sensitivity and selectivity scores for each class, to measure the impact of our mitigation efforts. Comparing these metrics before and after mitigation measures provides insights into changes in model behaviour. Additionally, we analyze error rates, particularly false positives and negatives, to assess improvements in the model’s predictive accuracy across different classes. For Vision Transformers, attention map analysis is used to verify if attention distribution is now more balanced across various classes and scenarios.
Model Re-calibration: This involves dynamically adjusting class weights within the loss function, guided by real-time performance metrics. This ensures the model stays optimized to any shifts in class representation or emerging imbalances. When new class samples or data become available, the model undergoes re-training to stay aligned with the evolving operational context. The last step further includes an iterative refinement, monitoring and updating bias mitigation strategies based on the latest performance assessments and re-calibrations.
IV Experimental Evaluation
IV-A Test Setup
Test Set: The overall dataset comprises approximately 1.4 million images, with annotations across several classes. Our focus is on the Pedestrian (149,921), Cyclist(17,060), and Motorcyclist (16,779) classes, as it provides a rich source for analyzing biases in object detection.
Preprocessing Steps: Data preprocessing involved normalization of image pixel values to the [0,1] range, aligning annotations to ensure consistency, and resizing images to standard dimensions suitable for input requirements of chosen models.
Model Architectures: The study uses SqueezeNet and ResNet18 from the CNN family alongside Vision Transformer: ViT. SqueezeNet, known for its compressed size and robustness in feature extraction, and ResNet18, recognized for its efficiency in learning from residual connections, are expected to provide insights into traditional CNN performance. ViT, representing the Transformer family, is included to evaluate the effectiveness of its attention-based mechanism in handling class imbalances.
IV-B Hyperparameter Optimization
Initial Hyperparameters: Were set as follows: learning rate of 0.001, batch size of 32, and weight decay of 0.0001 for CNN models. For ViT we configured the training with a batch size of 32 and an initial learning rate of 1e-3, which follows a linear decay to 1e-5. The model was trained for 30 epochs using the Adam optimizer. A dropout rate of 0.1 and a weight decay of 0.03 were applied to avoid overfitting.
Optimization Techniques: Hyperparameter tuning is performed using a grid search approach, systematically varying learning rates, batch sizes, and dropout rates to identify the optimal combination [2]. The optimal set of hyperparameters was chosen based on the highest average scores across these metrics as followed in [4], focusing on improvements in detecting classes.
IV-C Training Process
Training Details: As our sample size is small, CNN models were trained for 50 epochs using an Adam optimizer. A learning rate scheduler was used to reduce the learning rate by 10% every 10 epochs. Regularization techniques such as dropout and data augmentation (rotations and flipping) were used to prevent overfitting [16].
Data Splits: The dataset is divided into 70% training, 15% validation, and 15% test splits, ensuring a diverse representation of scenarios in each set.
Computational Resources: Training is conducted on a high-performance computing cluster with a combination of NVIDIA Tesla V100 GPUs using Open MPI and PyTorch as the deep learning framework.
IV-D Model Evaluation
Performance Metrics: Models were evaluated based on mAP, %IoU and NDS (Intersection over Union, NuScenes Detection Score), alongside sensitivity and selectivity scores for class-specific analysis.
Baseline Comparison: Model performances pre and post-implementation of bias mitigation strategies were compared. Baseline models were trained with respect to data conditions.
Class-specific Analysis: Performance metrics for each class were analyzed to assess improvements in detecting pedestrians, cyclists, and motorcyclists. Attention was paid to false positive or negative rate changes for these classes.
IV-E Bias Impact Analysis
Error Analysis: Figure 2 shows error analysis for the discussed classes. The ResNet18 model exhibited higher false positive rates across all classes than the ViT model, indicating a tendency towards over-prediction. Conversely, the ViT model maintained lower false positives but demonstrated higher false negatives, particularly for Pedestrians and Cyclists, suggesting a cautious but potentially under-detecting approach. Notably, the ViT model was more balanced for the Motorcyclist class, with significantly lower false negatives than ResNet18, indicating more reliable detection.
Visualizations: Figure 3 shows heatmap visualization to illustrate the models’ focus areas and any changes before the mitigation strategies. When trained with unbalanced classes, the tests show that both CNN models falsely classify motorcyclist and cyclist classes.
Figure 2: Figure shows class-wise error rate for ResNet and ViT according to False Negatives and False Positives in model validation
IV-F Discussion of Results
Figure 3: Figure showing heatmap for all three classes
TABLE I: NDS and %IoU for classes with ResNet18
%IoU of Classes NDS Data Condition
Pedestrian Cyclist Motorcyclist
86.1 53.6 76.4 72.0 Normal
78.5 8.3 40.7 42.5 Night
67.3 7.5 18.9 31.2 Weather
85.7 52.8 76.1 71.5 Rotated
72.1 12.4 21.4 35.3 Mixed
For measuring IoU, mAP and NDS scores mentioned in Tables 1 and 2, we used a subset approach, where the representation of one class (e.g., pedestrian) is kept at 67% and the other two classes have equal representation. This ratio is adjusted during iterations until all classes have the same representation. The purpose is to capture metrics in a scenario where the weighted class functions are normalized. Table I shows the Intersection over Union (%IoU) for different classes and the NuScenes Detection Score (NDS) across various conditions for ResNet18. Pedestrian detection remains relatively high in all conditions, with %IoU above 70% except when combined adverse conditions are present. Cyclist detection suffers significantly in all but normal conditions, with %IoU dropping to as low as 7.5% during adverse weather. Motorcyclist detection is also affected by these conditions but to a lesser extent. The overall NDS reflects these trends, with the highest score of 72.0 in normal conditions and the lowest of 31.2 during adverse weather. The rotated condition only slightly impacts the %IoU for pedestrians and motorcyclists but more so for cyclists. Mixed conditions present a challenging scenario with a noticeable decrease in %IoU across all classes and a resultant NDS of 35.3.
TABLE II: Mean Average Precision (mAP) of different classes With SqueezeNet under various conditions.
Pedestrian Cyclist Motorcyclist Total Data Condition
86.3 54.7 77.8 72.93 Normal
78.1 7.5 38.3 41.3 Night
67.0 8.3 19.7 31.6 Weather
85.9 52.3 75.2 71.1 Rotated
73.2 14.8 22.1 36.7 Mixed
Table II shows the mean average precision (mAP) of pedestrian, cyclist, and motorcyclist detection under various conditions. Under normal conditions, pedestrian detection is high at 86.3%, while cyclist detection lags at 54.7%. Motorcyclist detection is reasonably high at 77.8%, leading to an overall mAP of 72.93%. However, performance drops significantly at night and during adverse weather, with the lowest cyclist mAP at 7.5% and 8.3%, respectively. Interestingly, rotating images in normal conditions slightly reduces pedestrian and motorcyclist detection by about 1%, but notably impacts cyclist detection, dropping to 52.3%. The combination of all adverse conditions drastically reduces overall performance to a mAP of 36.7%, showcasing the challenges faced by the models in less-than-ideal scenarios.
TABLE III: Class-wise Performance Metrics on the CNN models
Class Model %IoU NDS Sens. Sel.
Pedestrian ResNet18 75.3 82.1% 0.78 0.80
Cyclist ResNet18 68.4 79.5% 0.65 0.67
Motorcyclist ResNet18 70.2 80.3% 0.70 0.72
Pedestrian SqueezeNet 78.1 85.4% 0.81 0.83
Cyclist SqueezeNet 72.6 82.7% 0.73 0.75
Motorcyclist SqueezeNet 74.0 83.2% 0.76 0.77
Sens. = Sensitivity Score, Sel = Selectivity Score
TABLE IV: Baseline vs Post-Mitigation Performance
Model Metric (Avg) Baseline Post Mitigation
ResNet18 %IoU 71.3% 75.6%
ViT %IoU 74.9% 79.2%
ResNet18 NDS 80.6% 83.7%
ViT NDS 83.8% 87.1%
Class-wise Performance Metrics: Table III shows class-wise performance metrics for two models, ResNet18 and SqueezeNet. For the pedestrian class, SqueezeNet outperforms ResNet18 with a higher %IoU and NDS. Cyclists and motorcyclists also see better %IoU and NDS scores with SqueezeNet. Sensitivity and selectivity scores, which measure the models’ ability to identify and differentiate classes, are consistently higher for SqueezeNet across all classes, indicating a more refined recognition capability. These metrics show overall better performance from SqueezeNet models in distinguishing and correctly identifying the classes.
Baseline vs Post-Mitigation Performance: Table IV shows average Intersection over Union (IoU) and NuScenes Detection Score (NDS) before and after applying bias mitigation techniques for ResNet18 and Vision Transformer (ViT). Both models show an improvement in %IoU and NDS after mitigation, with ViT exhibiting a larger increase in NDS. ResNet18’s IoU improved by 4.3%, and its NDS by 3.1%, while ViT’s IoU and NDS improved by 4.3% and 3.3%, respectively. This demonstrates the effectiveness of the bias mitigation strategies.
IV-G Key Takeaways
  • For CNN models, error analysis shows ResNet18 has higher false positive rates than SqueezeNet.
  • Post-mitigation performance, which involves resampling approaches, shows improvements in both CNN (ResNet18) and ViT models.
  • Under all conditions, the pedestrian class had higher detection. However, overall performance metrics show a noticeable decrease, seeking robust model development.
  • Equal number of samples from classes in a subset, does not necessarily ensure an equal representation of behaviour metrics during model learning.
V Conclusion and Future Work
This study explored the critical issue of class imbalance in driving dataset, specifically targeting the under-represented classes and its impact on the performance of AI models. Through a comprehensive methodology encompassing dataset analysis, model evaluation, and bias impact assessment, we detect disparities in the representation and detection of vulnerable road users. Our tests with popular CNN models and Vision Transformers have revealed how dataset biases can lead to alternate learning outcomes, adversely affecting the accuracy and correctness of perception systems. Further, we have implemented and evaluated bias mitigation techniques, which include cost-sensitive learning and targeted data augmentation. These methods have shown promising results in improving model performance, especially in IoU and NDS metrics. Furthermore, introducing a dynamic framework for ongoing bias assessment and model recalibration signifies an important step towards developing more equitable AI systems in autonomous driving.